OpenStudy (anonymous):

Solve: 1 over 2 - x= x over 6

OpenStudy (anonymous):

\[\frac{1}{2-x}=\frac{x}{6}\] this?

OpenStudy (anonymous):



OpenStudy (anonymous):

that

OpenStudy (anonymous):

make same denominator

OpenStudy (anonymous):

which would be 6 right

OpenStudy (nubeer):